OpenStudy (anonymous):

solve by elimination 3x-4y=-10 2x+5y=70

jhonyy9 (jhonyy9):

multiply first by 2 and second by -3 and assum 6x-8y=-20 -6x-15y=-210 --------------- 0 -23y= - 230 y= -230/(-23) = 10 y=10 so than 3x-4(10)= -10 so 3x-40= -10 so 3x= 30 so x= 30/3 =10